if the initial pressure is 1.01x10^5Pa and the initial temperature is 27°C, how much work does the gas in adiabatic process do during the compression if the initial volume of the cylinder is 1L? the compression ratio of a diesel engine is 15 to 1: that is, air in a cylinder is compressed to 1/15 of its original volume. use the values Cv=20.8J/mol.K and y=1.4 for air.


1
Expert's answer
2021-03-14T19:20:22-0400
"W=\frac{P_1V_1}{1-\gamma}\left(\left(\frac{V_2}{V_1}\right)^{1-\gamma}-1\right)\\
W=\frac{(1.01\cdot10^{5})(10^{-3})}{1-1.4}\left(\left(\frac{1}{15}\right)^{1-1.4}-1\right)=-493.43\ J"